I need to know how to solve these equations using substitution thanks:)
-x+y=7 and -1/2x-y=2

Answer by stanbon(75887)

solve these equations using substitution thanks:)
-x+y=7
-1/2x-y=2
===============
Solve either of the equations for x or for "y":
y = x+7
---
Substitute into the other equation:
(-1/2)x -(x+7) = 2
(-1/2)x - x - 7 = 2
---
(-3/2)x = 9
x = (-2/3)9
x = -6
-----
solve for "y":
y = x + 7
y = -6+7
y = 1
=============
Cheers,
Stan H.